OpenStudy (pphalke):

In how many ways can three committee members be chosen from a pool of 15 candidates?

OpenStudy (welshfella):

15C3 the number of combinations of 3 in 15 do you know the formula for this?

OpenStudy (pphalke):

no

OpenStudy (welshfella):

nCr = n! ------ r! (n-r)!

OpenStudy (pphalke):

ok so what do i do next

OpenStudy (welshfella):

plug in n = 15 and r = 3

OpenStudy (rolldextro):

455

OpenStudy (welshfella):

do you know what 15! ( 15 factorial) is?

OpenStudy (welshfella):

@rolldextro No direct answers please!

OpenStudy (pphalke):

no i dont

OpenStudy (rolldextro):

I am typing the solution..

OpenStudy (welshfella):

15! = 15*14*13.... * 1

OpenStudy (welshfella):

A quick way of doing this is to just go down 3 steps from 15 - that is 15*14*13 then divide by 3! which is 3*2*1 so we have 15C3 = 15*14*13 -------- 3*2*1 work that out and you have the answer.